OpenStudy (anonymous):

@amonaroll32 A credit card issuer charges an APR of 13.99%, and its billing cycle is 30 days long. What is its periodic interest rate? A. 14.37% B. 1.15% C. 0.97% D. 11.43%

OpenStudy (anonymous):

i think there is a formula i need to use but i dont know it

OpenStudy (anonymous):

Sorry, I'm not sure what this one would be. It'd be nice if you could stop mentioning my name though <.<

OpenStudy (tkhunny):

Well, A and D are just silly. Can you hazard a guess at which of the remaining two might be a better choice? Keep in mind that 13.99 > 12.

OpenStudy (anonymous):

b

OpenStudy (tkhunny):

Let's see... \(1.1399^{30/360} = 1.01097 \implies 1.097\)% \(1.1399^{30/365} = 1.01082 \implies 1.082\)% It's not real clear exactly how it is being calculated, but I would have to take a shot at B.

OpenStudy (tkhunny):

Just for the record: \(1.1399^{n/365} = 1.0115 \implies n = 31.8736\) \(1.1399^{n/360} = 1.0115 \implies n = 31.4370\) So, I'm still not clear on exactly how ti was calculated. Those are pretty weird numbers of days.